OpenStudy (anonymous):

Subtract 3 from 9. Then, add 6 Which of the following expressions matches the statement above? A. 3 - 9 + 6 B. 9 - 3 - 6 C. 9 - 6 + 3 D. 9 - 3 + 6 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(9-3)+6

OpenStudy (anonymous):

9-3+6=6+6=12

OpenStudy (alexandervonhumboldt2):

D

OpenStudy (anonymous):

9-3+6 the answer is B

OpenStudy (catlover5925):

@dtuniyants it is D, B subtracts 6

OpenStudy (anonymous):

oops sorry my bad